# FoREST: Frame of Reference Evaluation in Spatial Reasoning Tasks

### v1

@@ 1 Introduction
Figure 1: Illustration of FoR classes. The cat is the locatum, the car is the relatum, and the arrow indicates the perspective.
@@
Our contribution111code and dataset available at GitHub repository. are summarized as follows, 1. We introduce the FoREST benchmark to systematically evaluate LLMs’ FoR comprehension in a QA setting. 2. We analyze the impact of FoR information on text-to-image generation using multiple diffusion models. 3. We propose a prompting approach that generates spatial information, which can be incorporated into QA and layout diffusion to enhance performance.
@@ 2 Spatial Primitives
Spatial Relations. When dealing with spatial knowledge representation and reasoning, three main relations categories are often considered, that is, directional, topological, and distance (Hernández, 1994; Cohn and Renz, 2008; Kordjamshidi et al., 2011). Directional describes an object’s direction based on specific coordinates. Examples of relations include left and right. Topological describes the containment between two objects, such as inside. Distance describes qualitative and quantitative relations between entities. Examples of qualitative are far, and quantitative are 3km.
@@
3. Internal Intrinsic describes a spatial relation from the relatum’s perspective, where the relatum contains the locatum. The bottom-right image in Figure 1 show this with the sentence, A cat is inside and back of the car from the car’s perspective.
4. Internal Relative describes a spatial relation from the observer’s perspective where the locatum is inside the relatum. The bottom-left image in Figure 1 show this FoR with the sentence, A cat is inside and to the left of the car from my perspective.
Figure 2: Pipeline for dataset creation, starting from selecting a locatum and relatum from available objects and then applying a spatial template to generate the spatial expression ($T$). FoRs are assigned based on the relatum’s properties. $T$ is then categorized based on the number of FoRs. For example, A cat is to the right of a dog (with two possible FoRs: external intrinsic and external relative) belongs to the A-split. Then, its disambiguated version (A cat is to the right of a dog from the dog’s perspective) is added to the C-split. Next, if applicable, a relatum’s orientation is included for visualization and question generation. Finally, Unity3D generates scene configurations, and question-answer pairs are created from $T$.
@@ 3 FoREST Dataset Construction
To systematically evaluate LLM on the frame of reference (FoR) recognition, we introduce the Frame of Reference Evaluation in Spatial Reasoning Tasks (FoREST) benchmark. Each instance in FoREST consists of a spatial context ($T$), a set of corresponding FoR ($FoR$) which is a subset of {external relative, external intrinsic, internal intrinsic, internal relative}, a set of questions and answers ($\{Q,A\}$), and a set of visualizations ($\{I\}$). An example of $T$ is A cat is to the right of a dog. A dog is facing toward the camera. The FoR of this expression is {external intrinsic, external relative}. A possible question-answer is $Q$ = Based on the camera’s perspective, where is the cat from the dog’s position?, $A$ = {left, right}. There is an ambiguity in the FoR for this expression. Thus, the answer will be left if the model assumes the external relative. Conversely, it will be right if the model assumes the external intrinsic. The visualization of this example is in Figure 2.
@@ 3.2 Categories based on Relatum Properties
Using the FoR classes in Section 2, we identified two key relatum properties contributing to FoR ambiguity. The first property is the relatum’s intrinsic direction. It creates ambiguity between intrinsic and relative FoR since spatial relations can originate from the relatum’s and observer’s perspectives. The second is the relatum’s affordance as a container. It introduces the ambiguity between internal and external FoR, as spatial relations can refer to the inside and outside of the relatum. Based on these properties, we define four distinct cases: Cow Case, Box Case, Car Case, and Pen Case.
Case 1: Cow Case. In this case, the selected relatum has intrinsic directions but does not have the affordance as the container for the locatum. An obvious example is a cow, which should not be a container but has a front and back. In such cases, the relatum potentially provides a perspective for spatial relations. The applicable FoR classes are $FoR$ = {external intrinsic, external relative}. We augment the C-split with expressions of this case but include the perspective to resolve their ambiguity. To specify the perspective, we use predefined templates for augmenting clauses, such as from {relatum}’s perspective for external intrinsic or from the camera’s perspective for external relative. For example, if the context is A cat is to the right of the cow, in the A-split. The counterparts included in the C-split are A cat is to the right of the cow from cow’s perspective. for external intrinsic and A cat is to the right of the cow from my perspective. for external intrinsic.
Case 2: Box Case. The relatum in this category has the property of being a container but lacks intrinsic directions, making the internal FoR applicable. An example is a box. The applicable FoR classes are $FoR$ = {external relative, internal relative}. To include their unambiguous counterparts in the C-split, we specify the topological relation to the expressions, $T$, by adding inside for internal relative and outside for external relative cases. For example, for the sentence A cat is to the right of the box., the unambiguous $T$ with internal relative FoR is A cat is inside and to the right of the box. The counterpart for external relative is A cat is outside and to the right of the box.
@@
Case 4: Pen Case. In this case, the relatum lacks both the intrinsic direction and the affordance as a container. An obvious example is a pen with neither left/right nor the ability to be a container. Lacking these two properties, the created context has only one applicable FoR, $FoR$ = { external relative}. Therefore, we can categorize this case into both splits without any modification. An example of such a context is The book is to the left of a pen.
@@ 3.3 Context Visualization
In our visualization, complexity arises when the relatum has an intrinsic direction within the intrinsic FoR, as its orientation can complicate the spatial representation. For example, for visualizing A cat is to the right of a dog from the dog’s perspective., the cat can be placed in different coordinates based on the dog’s orientation. To address this issue, we add a template sentence for each direction, such as <relatum> is facing toward the camera, to specify the relatum’s orientation of all applicable $T$ for visualization and QA. For instance, A cat is to the left of a dog. becomes A cat is to the left of a dog. The dog is facing toward the camera. To avoid occlusion issues, we generate visualizations only for external FoRs, as one object may become invisible in internal FoR classes. We use only expressions in C-split since those have a unique FoR interpretation for visualization. We then create a scene configuration by applying a predefined template, as illustrated in Figure 2. Images are generated using the Unity 3D simulator Juliani et al. (2020), producing four variations per expression $T$ with different backgrounds and object positions. Further details on the simulation process are in Appendix B.
@@ 3.4 Question-Answering Generation
We generate questions for all generated spatial expressions ($T$). Note that we include the relatum orientation for cases where the relatum has an intrinsic direction, as mentioned in the visualization. Our benchmark includes two types of questions. The first type asks for the spatial relation between two given objects from the camera’s perspective, following predefined templates such as, Based on the camera’s perspective, where is the locatum relative to the relatum’s position? Template variations are made based on GPT4o. The second type of question queries the spatial relation from the relatum’s perspective. This question type follows the same templates but replaces the camera with the relatum. The first type of question is generated for all $T$, while the second type is only generated for $T$ where the relatum has intrinsic direction and a perspective can be defined accordingly. Question templates are provided in Appendix B.3. Answers are determined based on the corresponding FoRs, the spatial relation in $T$, and the relatum’s orientation when applicable.
@@ 4 Models and Tasks
The FoREST benchmark supports multiple tasks, including FoR identification, Question Answering (QA) that requires FoR comprehension, and Text-to-Image (T2I). This paper focuses on QA and T2I for a deeper evaluation of spatial reasoning. FoR identification experiments are provided in Appendix E.
@@ 4.1 Question-Answering (QA)
Task. This QA task evaluates LLMs’ ability to adapt contextual perspectives across different FoRs. Both A and C splits are used in this task. The input is the context, consisting of a spatial expression $T$ and relatum orientation, if available, and a question $Q$ that queries the spatial relation from either an observer or the relatum’s perspective. The output is a spatial relation $S$, restricted to {left, right, front, back}.
Zero-shot baseline. We call the LLM with instructions, a spatial context, $T$, and a question, $Q$, expecting a spatial relation as the response. The prompt instructs the model to answer the question with one of the candidate spatial relations without any explanations.
Few-shot baseline. We create four spatial expressions, each assigned to a single FoR class to prevent bias. Following the steps in Section 3.4, we generate a corresponding question and answer for each. These serve as examples in our few-shot prompting. The input to the model is instruction, example, spatial context, and the question.
Chain-of-Thought baseline (Wei et al., 2023). To create Chain-of-Thought (CoT) examples, we modify the prompt to require reasoning before answering. We manually crafted reasoning explanations with the necessary information for each example we used in the few-shot setting. The input to the model is instruction, CoT example, spatial context, and the question.
@@
Table 1: QA accuracy in the A-Split across various LLMs. Rand Irepresent the percentage the model assumes relative or intrinsic FoR for ambiguous expression explained in Section 5.1. Acc is the accuracy, and Avg is the micro-average of accuracy. (1): 0-shot, (2): 4-shot, (3): CoT, and (4): SG + CoT.
@@ 4.2 Text-To-Image (T2I)
Task. This task aims to determine the diffusion models’ ability to consider FoR by evaluating their generated images. The input is a spatial expression, $T$, and the output is a generated image ($I$). We use the context from both C and A splits with external FoRs for this task.
@@
Layout Diffusion Models. We evaluate the Layout Diffusion model, a more advanced T2I model operating in two phases: text-to-layout and layout-to-image. Given that LLMs can generate the bounding box layout (Cho et al., 2023b), we provide them with instructions and $T$ to create the layout. The layout consists of bounding box coordinates for each object in the format of {object: $[x,y,w,h]$}, where $x$ and $y$ denote the starting point and $h$ and $w$ denote the height and width. The bounding box coordinates and $T$ are then passed to the layout-to-image model to produce the final image, $I$.
4.3 Spatial-Guide Prompting
We hypothesize that the spatial relation types and FoR classes explained in Section 2 can improve question-answering and layout generation. For instance, the external intrinsic FoR emphasizes that spatial relations originate from the relatum’s perspective. To leverage this, we propose Spatial-Guided (SG) prompting, an additional step applied before QA or layout generation steps. This step extracts spatial information, including direction, topology, distance as well as the FoR from spatial expression $T$. The extracted information will serve as supplementary for guiding LLMs in QA and layout generation. We manually craft four examples covering these aspects. First, we specify the perspective for directional relations, e.g., left relative to the observer, to distinguish intrinsic from relative FoR. Next, we indicate whether the locatum is inside or outside the relatum for topological relations to differentiate internal from external FoR. Lastly, we provide an estimated quantitative distance to support topological and directional relation identification, such as far. These examples are then provided as a few-shot example for the model to extracted information automatically.
@@ 5 Experimental Results
Qwen2-72B (SG + CoT) $97.6$ $42.5$ $31.3$ $93.8$ $71.4$ $42.8$ $86.6$ $92.0$ $34.0$ $64.5$
@@
GPT-4o (CoT) $81.7$ $\mathbf{76.1}$ $\mathbf{82.4}$ $71.5$ $78.8$ $53.0$ $91.1$ $90.6$ $\mathbf{50.8}$ $71.9$
GPT-4o (SG + CoT) $\mathbf{97.9}$ $72.2$ $72.7$ $\mathbf{93.4}$ $\mathbf{85.8}$ $48.9$ $96.3$ $95.9$ $36.1$ $\mathbf{71.8}$
@@
5.1 Evaluation Metrics
QA. We report an accuracy measure defined as follows. Since the questions can have multiple correct answers, specifically in A-split, as explained in Section 3, the prediction is correct if it matches any valid answer. Additionally, we report the model’s bias distribution when FoR ambiguity exists. $I$is the percentage of correct answers when assuming an intrinsic FoR, while $R$is this percentage with a relative FoR assumption. Note that cases where both FoR assumptions lead to the same answer are excluded from these calculations.
T2I. We adopt spatialEval (Cho et al., 2023b) approach for evaluating T2I spatial ability. However, we modify it to account for FoR. We convert all relations to a camera perspective before passing them to spatialEval, which assumes this viewpoint. Accuracy is determined by comparing the bounding box and depth map of the relatum and locatum. For FoR ambiguity, a generated image is correct if it aligns with at least one valid FoR interpretation. We report results using VISORcond and VISORuncond (Gokhale et al., 2023), metrics for assessing T2I spatial understanding. VISORcond evaluates spatial relations only when both objects appear correctly, aligning with our focus on spatial reasoning rather than object creation. In contrast, VISORuncond evaluates the overall performance, including object creation errors.
5.2 Experimental Setting
QA. We use Llama3-70B (Llama, 2024), Qwen2-72B (Qwen Team, 2024), and GPT-4o (gpt-4o-2024-11-20(OpenAI, 2024) as the backbones for prompt engineering. To ensure reproducibility, we set the temperature of all models to 0. For all models, we apply zero-shot, few-shot, CoT, and our proposed prompting with CoT (SG+CoT).
T2I. We select Stable Diffusion SD-1.5 and SD-2.1 (Rombach et al., 2021) as our stable diffusion models and GLIGEN(Li et al., 2023) as the layout-to-image backbone. For translating spatial descriptions into textual bounding box information, we use Llama3-8B and Llama3-70B, as detailed in Section 4.2. The same LLMs are used to generate spatial information for SG prompting. We generate four images to compute the VISOR score following Gokhale et al. (2023) Inference steps for all T2I models are set to 50. For the evaluation modules, we select grounding DINO (Liu et al., 2025) for object detection and DPT (Ranftl et al., 2021) for depth mapping, following VPEval Cho et al. (2023b). The experiments were conducted on an $A6000$ GPU, totaling approximately $300$ GPU hours.
5.3 Results
RQ1. What is the bias of the LLMs for the ambiguous FoR? Table 1 presents the QA results for the A-split. Ideally, a model that correctly extracts the spatial relation without considering perspective should achieve 100% accuracy, as the context lacks a fixed perspective. However, this ideal model is not the focus of our work. We aim to assess model bias by measuring how often LLMs adopt a specific perspective when answering. In the Cow and Pen case, relatum properties do not introduce FoR ambiguity in directional relations, making the task pure extraction rather than reasoning. Thus, we focus on the $I$and $R$of the Cow and Car cases, which best reflect LLMs’ bias. Qwen2 achieves around 80% accuracy across all experiments by selecting spatial expressions directly from the context, suggesting it may disregard the question’s perspective. GPT-4o shows similar bias in 0-shot and 4-shot settings but shifts toward intrinsic interpretation with CoT. This bias reduces accuracy in camera-perspective questions from 93.2% to 81.4%, where FoR adaptation is more challenging than relation extraction. Llama3-70B lacks a strong preference, balancing assumptions but slightly favoring relative FoR. This uncertainty lowers performance, requiring more reasoning to reach the correct answer. In summary, Qwen2 achieves higher accuracy by focusing on relation extraction without considering FoR reasoning, while other models attempt reasoning but struggle to reach correct conclusions, leading to lower performance.
@@ 5.3 Results
Table 3: VISORcond score on the A and C splits where $I$ refer to the Cow case and Car case where relatum has intrinsic directions, and $R$ refer to the Box case and Pen case where relatum lacks intrinsic directions, $avg$ is mirco-average of $I$ and $R$. cond are explained in Section 5.1. EI and ER FoR represent the generated image considered corrected by EI or ER FoR

### v2

@@ 1 Introduction
Figure 1: Illustration of FoR classes. The cat is the locatum, the car is the relatum, and the arrow denotes the perspective.
@@
Our contribution111code and dataset available at GitHub repository. are summarized as follows, 1. We introduce the FoREST benchmark to evaluate LLMs’ FoR comprehension in a QA setting systematically. 2. We analyze the impact of FoR information on text-to-image generation using multiple diffusion models. 3. We propose a prompting approach that generates spatial information, which can be incorporated into QA and layout diffusion to enhance performance.
@@ 2 Spatial Primitives
Spatial Relations. When dealing with spatial knowledge representation and reasoning, three main relations are often considered, directional, topological, and distance (Hernández, 1994; Cohn and Renz, 2008; Kordjamshidi et al., 2011). Directional describes an object’s direction based on specific coordinates, e.g., left and right. Topological describes the containment between two objects, such as inside. Distance describes qualitative and quantitative relations between entities. Examples of qualitative are far, and quantitative are 3km.
@@
3. Internal Intrinsic describes a spatial relation from the relatum’s perspective, where the relatum contains the locatum. The bottom-right image in Figure 1 shows this with the sentence, A cat is inside and back of the car from the car’s perspective.
4. Internal Relative describes a spatial relation from the observer’s perspective where the locatum is inside the relatum. The bottom-left image in Figure 1 shows this FoR with the sentence, A cat is inside and to the left of the car from my perspective.
Figure 2: Pipeline for dataset creation, starting from selecting a locatum and relatum from available objects and then applying a spatial template to generate the spatial expression ($T$). FoRs are assigned based on relatum properties. $T$ is categorized based on the number of FoRs. For example, A cat is to the right of a dog (with two possible FoRs: external intrinsic and external relative) belongs to the A-split. Then, its disambiguated version (A cat is to the right of a dog from the dog’s perspective) is added to the C-split. Next, if applicable, a relatum’s orientation is included for visualization and question generation. Finally, Unity3D generates scene configurations, and question-answer pairs are derived from $T$.
@@ 3 FoREST Dataset Construction
To systematically evaluate LLM on the frame of reference (FoR) recognition, we introduce the Frame of Reference Evaluation in Spatial Reasoning Tasks (FoREST) benchmark. Each instance in FoREST consists of a spatial context ($T$), a set of corresponding FoR ($FoR$) which is a subset of {external relative, external intrinsic, internal intrinsic, internal relative}, a set of questions and answers ($\{Q,A\}$), and a set of visualizations ($\{I\}$). An example of $T$ is A cat is to the right of a dog. A dog is facing toward the camera. The FoR of $T$ is {external intrinsic, external relative}. A question-answer is $Q$ = Based on the camera’s perspective, where is the cat from the dog’s position?, $A$ = {left, right}. There is an ambiguity in the FoR for this expression. Thus, the answer will be left if the model assumes the external relative. Conversely, it will be right if the model assumes the external intrinsic. The visualization of this example is in Figure 2.
@@ 3.2 Categories based on Relatum Properties
Using the FoR classes in Section 2, we identify two key properties contributing to FoR ambiguity. The first property is the relatum’s intrinsic direction. It creates ambiguity between intrinsic and relative FoRs as spatial relations may originate from either relatum’s or observer’s perspectives. The second is the relatum’s affordance as a container. It introduces the ambiguity between internal and external FoR, as spatial relations may refer to either inside or outside of the relatum. Note that containment is defined as the ability of the relatum to contain the locatum, considering both objects’ sizes. Based on these properties, we define four distinct cases: Cow, Box, Car, and Pen.
Case 1: Cow Case. In this case, the selected relatum has intrinsic directions but cannot be the container for the locatum. An example is a cow. In such cases, the relatum potentially provides a perspective for spatial relations. The applicable FoR classes are $FoR$ = {external intrinsic, external relative}. We augment the C-split with expressions of this case but include the perspective to resolve their ambiguity. To specify the perspective, we use predefined templates for augmenting clauses, such as from {relatum}’s perspective for external intrinsic or from the camera’s perspective for external relative. For example, if the context is A cat is to the right of the cow, in the A-split. The counterparts included in the C-split are A cat is to the right of the cow from cow’s perspective. for external intrinsic and A cat is to the right of the cow from my perspective. for external intrinsic.
Case 2: Box Case. The relatum in this case has the property of being a container but lacks intrinsic directions, making the internal FoR applicable. An example is a box. The applicable FoR classes are $FoR$ = {external relative, internal relative}. To include their unambiguous counterparts in the C-split, we specify the topological relation by adding inside for internal relative and outside for external relative. For example, for the sentence A cat is to the right of the box., the unambiguous $T$ with internal relative FoR is A cat is inside and to the right of the box. The counterpart for external relative is A cat is outside and to the right of the box.
@@
Case 4: Pen Case. In this case, the relatum lacks both the intrinsic direction and the affordance as a container. An obvious example is a pen. Lacking these two properties, the created context has only one applicable FoR, $FoR$ = {external relative}. Therefore, we can categorize this case into both splits without any modification. An example of such a context is The book is to the left of a pen.
@@ 3.3 Context Visualization
In our visualization, complexity arises when the relatum has an intrinsic direction within the intrinsic FoR, as its orientation can complicate the spatial representation. For example, for visualization of A cat is to the right of a dog from the dog’s view, the cat can be placed in different coordinates based on the dog’s orientation. To address this issue, we add a template sentence for each direction, such as <relatum> is facing toward the camera, to specify the relatum’s orientation of all applicable $T$ for visualization and QA. For instance, A cat is to the left of a dog. becomes A cat is to the left of a dog. The dog is facing toward the camera. To avoid occlusion issues, we generate visualizations only for external FoRs, as one object may become invisible in internal FoR classes. We use only expressions in C-split since those have a unique FoR interpretation for visualization. We then create a scene configuration by applying a predefined template, as illustrated in Figure 2. Images are generated using the Unity 3D simulator Juliani et al. (2020), producing four variations per expression $T$ with different backgrounds and object positions. Further details on the simulation process are in Appendix A.2.
@@ 3.4 Question-Answering Generation
We generate questions for all generated spatial expressions ($T$). Note that we include the relatum orientation for cases where the relatum has an intrinsic direction, as mentioned in the visualization. Our benchmark includes two types of questions. The first type asks for the spatial relation between two given objects from the camera’s perspective, following predefined templates such as, Based on the camera’s perspective, where is the locatum relative to the relatum’s position? Template variations are made based on GPT4o; details are provided in Appendix A.2.4. The second type of question queries the spatial relation from the relatum’s perspective. This question type follows the same templates but replaces the camera with the relatum. The first type of question is generated for all $T$, while the second type is only generated for $T$ where the relatum has intrinsic direction and a perspective can be defined accordingly. Answers are determined based on the corresponding FoRs, the spatial relation in $T$, and the relatum’s orientation when applicable.
@@ 4 Models and Tasks
This paper focuses on Question-Answering and Text-to-Image using the FoREST benchmark to evaluate FoR in spatial reasoning comprehensively. FoREST also supports other tasks, such as FoR identification, detailed in Appendix H.
@@ 4.1 Question-Answering (QA)
Task. This QA task evaluates LLMs’ ability to adapt contextual perspectives across different FoRs. The input includes a spatial expression $T$, relatum orientation (if available), and a question $Q$ querying the spatial relation from either the observer’s or the relatum’s perspective. The output is a spatial relation $S$, restricted to {left, right, front, back}. We use the context from both splits for this task.
Zero-shot baseline. We call the LLM with instructions, a spatial context, and a question expecting a spatial relation as the response. The prompt instructs the model to answer the question with one of the candidate spatial relations without explanation.
Few-shot baseline. We create four spatial expressions, each assigned to a single FoR class to prevent bias. Following the steps in Section 3.4, we generate a corresponding question and answer for each. These serve as examples in our few-shot prompting. The input to the model is the instruction, example, spatial context, and the question.
Chain-of-Thought (CoT) baseline (Wei et al., 2023). To create CoT examples, we modify the prompt to require reasoning before answering. We manually craft reasoning explanations with the necessary information for the few-shot examples. The input to the model is the instruction, CoT example, spatial context, and the question.
Human baseline. We conducted a small scope human study with three participants, paid as research assistants, to establish a human baseline for A-split. Each participant is shown 25 randomly sampled QA examples per relatum case, totaling 150 examples. Each includes a textual scene description and corresponding question. See Appendix I for details. We report average accuracy across participants to reflect the human baseline.
@@
Qwen2VL-72B (1) $49.5$ $\mathbf{50.5}$ $78.3$ $\mathbf{51.6}$ $48.4$ $80.0$ $98.3$ $96.9$ $81.2$ $41.0$ $\mathbf{59.0}$ $55.4$ $44.7$ $\mathbf{55.3}$ $59.0$ $56.0$
Qwen2VL-72B (2) $40.9$ $\mathbf{59.1}$ $89.4$ $44.7$ $\mathbf{55.3}$ $79.0$ $100.0$ $100.0$ $89.6$ $32.7$ $\mathbf{67.3}$ $66.9$ $28.5$ $\mathbf{71.5}$ $67.0$ $66.9$
Qwen2VL-72B (3) $\mathbf{63.5}$ $36.5$ $84.0$ $\mathbf{72.2}$ $27.8$ $84.5$ $100.0$ $100.0$ $86.3$ $\mathbf{51.0}$ $49.0$ $77.7$ $\mathbf{54.6}$ $45.4$ $82.8$ $78.4$
Qwen2VL-72B (4) $\mathbf{50.5}$ $49.5$ $78.3$ $\mathbf{59.5}$ $40.5$ $67.4$ $98.3$ $99.6$ $79.9$ $23.5$ $\mathbf{76.5}$ $78.2$ $33.0$ $\mathbf{67.0}$ $67.7$ $76.7$
@@
Human-baseline $36.6$ $\mathbf{63.4}$ $90.7$ $27.8$ $\mathbf{72.2}$ $96.0$ $72.0$ $82.7$ $85.3$ $41.4$ $\mathbf{58.6}$ $97.3$ $36.1$ $\mathbf{63.9}$ $96.0$ $96.7$
Table 1: QA accuracy in the A-Split. Rand Irepresent the percentage the model assumes relative or intrinsic FoR for ambiguous expression, explained in Section 5.1. Acc is the accuracy, and Avg is the micro-average of accuracy. (1): 0-shot, (2): 4-shot, (3): CoT, and (4): SG+CoT.
@@ 4.2 Text-To-Image (T2I)
Task. This task assesses the diffusion models’ ability to consider FoR by evaluating their generated images. The input is a spatial expression, $T$, and the output is a generated image ($I$). C and A splits with external FoRs are used for this task.
@@
Layout Diffusion Models. This model operates in two phases: text-to-layout and layout-to-image. Given that LLMs can generate the bounding box layout (Cho et al., 2023b), we provide them with instructions and $T$ to create the layout. The layout consists of bounding box coordinates for each object in the format of {object: $[x,y,w,h]$}, where $x$ and $y$ denote the starting point and $h$ and $w$ denote the height and width. The bounding box coordinates and $T$ are then passed to the layout-to-image model to produce the final image, $I$.
4.3 Spatial-Guide (SG) Prompting
We hypothesize that the spatial relation types and FoR classes defined in Section 2 can improve QA and layout generation. For example, the external intrinsic FoR emphasizes that spatial relations originate from the relatum’s perspective. To leverage this, we propose SG prompting, an additional step applied before QA or layout generation. This step extracts spatial information, including direction, topology, distance, and the FoR, from the spatial expression $T$, as supplementary input to guide LLMs in QA or layout generation. We manually craft four examples covering these aspects. First, we specify the perspective for directional relations, e.g., left relative to the observer, to distinguish intrinsic from relative FoR. Next, we indicate whether the locatum is inside or outside the relatum for topological relations to differentiate internal from external FoR. Lastly, we provide an estimated quantitative distance to support topological and directional relation identification, e.g., far. These examples are provided in a few-shot setting to guide the model in automatically extracting such information.
@@ 5 Experimental Results
Qwen2-72B (SG+CoT) $97.6$ $42.5$ $31.3$ $\mathbf{93.8}$ $71.4$ $42.8$ $86.6$ $92.0$ $34.0$ $64.5$
Qwen2VL-72B (0-shot) $68.3$ $40.2$ $42.4$ $75.3$ $56.0$ $37.4$ $60.1$ $82.2$ $35.5$ $50.0$
Qwen2VL-72B (4-shot) $78.4$ $41.0$ $44.5$ $66.1$ $60.5$ $38.0$ $85.4$ $91.2$ $29.9$ $61.6$
Qwen2VL-72B (CoT) $58.4$ $62.4$ $68.4$ $40.7$ $59.3$ $48.6$ $43.5$ $36.1$ $\mathbf{61.9}$ $46.4$
Qwen2VL-72B (SG+CoT) $\mathbf{99.0}$ $50.0$ $47.5$ $93.0$ $75.9$ $27.3$ $37.3$ $37.1$ $34.0$ $32.7$
@@
GPT-4o (CoT) $81.7$ $\mathbf{76.1}$ $\mathbf{82.4}$ $71.5$ $78.8$ $53.0$ $91.1$ $90.6$ $50.8$ $\mathbf{71.9}$
GPT-4o (SG+CoT) $97.9$ $72.2$ $72.7$ $93.4$ $\mathbf{85.8}$ $48.9$ $96.3$ $95.9$ $36.1$ $71.8$
@@
Table 3: VISORcond score explained in Section 5.1 where $I$ refers to the Cow and Car cases where relatum has intrinsic directions, and $R$ refers to the Box and Pen cases where relatum lacks intrinsic directions, $avg$ is the micro-average of $I$ and $R$. EI and ER FoR represent the generated image considered corrected by EI or ER FoR.
5.1 Evaluation Metrics
QA. We report an accuracy (acc.) based on the correct answer defined as follows. Since the questions can have multiple correct answers, as explained in Section 3, the prediction is correct if it matches any valid answer. Additionally, we report the model’s bias distribution when FoR ambiguity exists. $I$is the percentage of correct answers when assuming an intrinsic FoR, while $R$is this percentage with a relative FoR assumption. Note that cases where both FoR assumptions lead to the same answer are excluded from the bias calculation.
T2I. We adopt spatialEval (Cho et al., 2023b) for evaluating T2I spatial ability. We modify it to account for FoR by converting relations to a camera perspective before passing them to spatialEval, which assumes this viewpoint. Accuracy is determined by comparing the bounding box and depth map of the relatum and locatum. For FoR ambiguity, a generated image is correct if it aligns with at least one valid FoR interpretation. We report results using VISORcond and VISORuncond (Gokhale et al., 2023) metrics. VISORcond evaluates spatial relations only when both objects appear correctly, aligning with our focus on spatial reasoning. While, VISORuncond evaluates the overall performance, including object creation errors.
5.2 Experimental Setting
QA. We use Llama3-70B (Llama, 2024), Qwen2-72B (Qwen Team, 2024), Qwen2VL-72B Wang et al. (2024), and GPT-4o (OpenAI, 2024) as backbone models. All models are evaluated with zero-shot, few-shot, CoT, and our SG+CoT prompting under temperature 0 to ensure reproducibility.
T2I. We select Stable Diffusion SD-1.5 and SD-2.1 (Rombach et al., 2021) as our stable diffusion models and GLIGEN(Li et al., 2023) as the layout-to-image backbone. For translating spatial descriptions into bounding box information, we use Llama3-8B and Llama3-70B, as detailed in Section 4.2. The same LLMs are used to generate spatial information for SG prompting. We generate four images to compute the VISOR score following Gokhale et al. (2023) Inference steps for all T2I models are set to 50. For the evaluation, we select grounding DINO (Liu et al., 2025) for object detection and DPT (Ranftl et al., 2021) for depth mapping, following VPEval Cho et al. (2023b). The experiments were conducted on two $A6000$ GPUs, totaling approximately $300$ GPU hours.
5.3 Results
RQ1. What is the LLM’s bias for the ambiguous FoR? Table 1 presents the QA results for the A-split. Ideally, a model extracting spatial relations without considering perspective should achieve 100% accuracy, as the context lacks a fixed perspective. However, our focus is not on this. We aim to assess LLMs’ bias by measuring how often they adopt a specific perspective when answering. In the Box and Pen cases, relatum properties do not introduce FoR ambiguity in directional relations, making the task pure extraction rather than reasoning. Thus, we focus on the $I$and $R$of the Cow and Car cases, which best reflect LLMs’ bias. Qwen2 achieves 80% accuracy across all experiments by selecting spatial relations directly from context, suggesting it may disregard the question’s perspective. This is supported by attention analysis in Appendix C. The attention maps show that most correct Qwen2 responses assume a shared perspective between context and question and pay low attention to the perspective mentioned in the question. GPT-4o exhibits similar bias in 0-shot and 4-shot settings but shifts toward intrinsic interpretation with CoT. This shift reduces accuracy in camera-perspective questions where FoR adaptation is more significant than relation extraction. Llama3-70B shows no strong preference, slightly favoring relative FoR, but this balanced assumption leads to lower performance due to increased reasoning demands. Qwen2VL exhibits a similar pattern but achieves higher accuracy, likely due to visual training. GPT-o4-mini with high visual reasoning shows the same pattern, which is later discussed as an additional experiment in Appendix J. Moreover, our experiments with humans, presumed to have prior visual knowledge, show strong reasoning ability, achieving over 90% accuracy even in scenarios where perspective preference requires FoR adaptation. In our results, humans slightly prefer intrinsic FoR over relative FoR. Their cultural background and recently seen examples may also influence their preference as discussed in Coventry et al. (2018). A more detailed analysis of human performance is provided in Appendix I. In summary, we conjecture that Qwen2 performs well by focusing on extraction without reasoning FoR, while other models attempt reasoning but often fail to reach correct conclusions, leading to lower accuracy.